Mass 1 is initially moving 7.5m/s in the +x direction and collides in a perfectly elastic ally with Mass 2 moving in the -x direction at 8.4 m/s. After the collision, Mass 1 is now moving in the -x direction at 6.7m/s. What is the velocity of Mass 2 after the collision in m/s? Indicate -x direction by including a negative sign.
Answer #1

In an elastic collision relative velocity of one mass in relative to other remains same in magnitude and opposite in direction.

Before collision,

velocity of mass 2 relative to mass 1 = -8.4 m/s - 7.5 m/s = -15.9 m/s

After collision,

velocity of mass 2 relative to mass 1 = 15.9 m/s

velcity of mass 1 = -6.7 m/s

hence, velocity of mass 2 = (15.9 - 6.7) m/s = 9.2 m/s